# Radar-Guided Polynomial Fitting for Metric Depth Estimation

### v1

@@ 3.1 Motivation
indicating where the second derivative changes sign. Our method learns inflection points to model transition regions where the curvature of the MDE error changes, allowing necessary non-uniform corrections.
@@
       50m Lin [26] [IROS ’20] 1598.2 3790.1 1576.3 3464.7 2894.9 5507.7
RadarNet [45] [CVPR ’23] 1727.7 3746.8 1430.5 3250.8 2944.8 6113.2
SparseBeatsDense [24] [ECCV ’24] 1524.5 3567.3 1424.4 3267.5 2909.9 5746.4
GET-UP [49] [WACV ’25] 1241.0 2857.0 1483.9 3220.5 2330.0 4565.0
RadarCam-Depth [25] [ICRA ’24] 1286.1 2964.3 1067.5 2817.4 1895.1 4458.7
PolyRad (Ours) 1146.7 2695.6 578.0 1108.6 1293.7 2420.6
       70m Lin [26] [IROS ’20] 1897.8 4558.7 1701.7 3853.3 3312.3 6480.5
RadarNet [45] [CVPR ’23] 2073.2 4590.7 1543.8 3655.3 3428.9 7331.9
SparseBeatsDense [24] [ECCV ’24] 1822.9 4303.6 1520.0 3593.4 3408.3 6914.7
GET-UP [49] [WACV ’25] 1541.0 3657.0 1651.5 3711.7 2758.0 5678.0
RadarCam-Depth [25] [ICRA ’24] 1587.9 3662.5 1157.0 3117.7 2095.4 4944.6
PolyRad (Ours) 1402.9 3269.7 603.7 1154.9 1442.9 2803.8
       80m Lin [26] [IROS ’20] 1988.4 4841.1 1739.5 3993.9 3468.1 6910.3
RadarNet [45] [CVPR ’23] 2179.3 4898.7 1578.4 3804.2 3597.0 7809.2
SparseBeatsDense [24] [ECCV ’24] 1927.0 4609.6 1548.4 3708.1 3584.0 7375.2
GET-UP [49] [WACV ’25] 1632.0 3932.0 1699.7 3882.6 2917.3 6145.1
RadarCam-Depth [25] [ICRA ’24] 1689.7 3948.0 1183.5 3229.0 2227.4 5385.8
PolyRad (Ours) 1454.1 3488.9 629.6 1171.3 1500.1 3951.8
Table 1: Quantitative results on nuScenes, ZJU-4DRadarCam, and View-of-Delft test sets evaluated with various maximum depths.
3.2 Radar Processing
We begin with a radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$, where $N_{C}$ denotes the number of points and each point is represented using its coordinates in three-dimensional camera space. Radar point clouds, while noisy and sparse, provide metric depth measurements. We encode $C$ using a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
To facilitate effective radar feature aggregation, we introduce a set of learnable prototypes $P\in\mathbb{R}^{N_{P}\times c_{r}}$, where these prototypes can be interpreted as capturing diverse spatial and geometric properties present in the radar point cloud. Each prototype learns to focus on specific patterns in the configuration of radar points, making them highly expressive and adaptable across varying scenes. Unlike existing works [45, 25] that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud. This allows for selective feature aggregation that mitigates the impact of outliers, resulting in more robust depth predictions.
The next step is to match the learnable prototypes to the most relevant radar features for a given input. We allow each prototype to dynamically attend to the radar points that provide the most relevant spatial and geometric signals, thereby aggregating a sparse point cloud into a global scene-descriptive representation. To achieve this, we project the point-wise radar features $F_{r}$ into a key-value space, where they can be efficiently retrieved by the prototypes $P$. We compute the aggregated radar features $R$:
$R=\text{softmax}\left(\frac{P\times(\psi_{r}^{K}(F_{r}))^{T}}{\sqrt{c_{r}}}\right)\psi_{r}^{V}(F_{r}),$ (3)
where $\psi_{r}^{K}$ and $\psi_{r}^{V}$ are MLPs that project the point-wise radar features $F_{r}$. Comparatively, applying depth completion methods that densify a sparse projection of points using surrounding context [31, 36] to a radar point cloud results in poor performance (see Supp. Mat). This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution [45].
Next, the scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the MDE model is encoded using a depth encoder, $f_{z}$, to produce depth features $Z\in\mathbb{R}^{(H\times W)\times c_{z}}$. We capitalize on the invariants learned through large-scale MDE training—such as robustness to color variations, illumination changes, and diverse object poses—by extracting features from the MDE depth maps, which leads to features encoding shapes and surfaces. We note that variations in the shape of objects or their geometry are comparatively smaller than variations in visual appearance, which are highly sensitive to transient factors such as perspective and lighting. We then learn spatial correspondences between the depth features $Z$ and the radar features $R$ to produce a unified scene representation that fuses the structural information encoded in the MDE predictions with the metric cues provided by the radar features. An important insight is that our depth encoder $f_{z}$ yields more stable geometric context by matching radar point configurations to shapes and surfaces that remain relatively consistent across scenes, rather than to pixel-wise color intensities that can vary arbitrarily with lighting and viewpoint.
To this end, we fuse the depth features $Z$ and the aggregated radar features $R$ to yield a unified scene representation
$S=\text{softmax}\left(\frac{Z\times(\psi_{R}^{K}(R))^{T}}{\sqrt{c_{r}}}\right)\psi_{R}^{V}(R),$ (4)
where $\psi_{R}^{K}$ and $\psi_{R}^{V}$ are MLPs that project the aggregated radar features $R$. The fused representation $S$ is passed through a shallow convolutional neural network (CNN) $f_{s}$ followed by a global average pooling (GAP) layer to yield
$\bar{S}=\text{GAP}(f_{s}(S)),$ (5)
which is a $c_{s}$-dimensional feature vector. The GAP layer aggregates spatial information from the entire scene, ensuring that $\bar{S}$ captures global context. This final scene representation $\bar{S}$ is then fed into an MLP $\psi_{s}$, which predicts the $N+1$ polynomial coefficients as a vector $\hat{c}$:
$\hat{c}=\psi_{s}(\bar{S})\in\mathbb{R}^{N+1}.$ (6)
These coefficients allow the model to adaptively refine the initial depth predictions from the MDE model, with each coefficient adjusting depth at different scales and granularities. Lower-order terms capture broad scene structure, while higher-order terms enable fine-grained and inter-region corrections, resulting in a high-fidelity metric-scale depth map.
Figure 3: Qualitative comparison on nuScenes. Baseline methods GET-UP and RadarCam-Depth (RC-D) fail to correctly reconstruct the 3D scene, resulting in objects with incorrect or missing depth predictions. While raw MDE outputs reasonable scaleless reconstructions, several cross-object misalignments are apparent. In contrast, PolyRad leverages polynomial fitting to infer high-fidelity depth predictions.
3.3 Polynomial Fitting
Given a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ predicted by the MDE foundation model, our goal is to transform it into a metric depth map $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ using the polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\dots,\hat{c}_{N}\}$ predicted by our network (Sec. 3.2).
Formally, we express the final depth map $\hat{d}$ as:
$\begin{split}\hat{d}(x,y)=\sum_{i=0}^{N}\,\hat{c}_{i}\cdot z(x,y)^{i},\;\;\forall(x,y)\in H\times W\end{split}$ (7)
This polynomial formulation applies the learned coefficients to progressively higher powers of the scaleless depth map, enabling complex, non-linear transformations that go beyond simple scale-and-shift adjustments. Each pixel’s final depth is thus computed by combining multiple weighted terms, where each term corresponds to a different polynomial order of the initial depth prediction.
This is in contrast to directly mapping to a metric-scale depth map [45, 25], where the degrees of freedom comprise every predicted pixel. Each corresponding point and pixel prediction can be viewed as a constraint or equation in our estimation system. As the radar point cloud (hundreds of points) only occupies a small subset of the image space, the solution is underdetermined and often irregular (see Fig. 3). Conversely, a simple linear fit (only two degrees of freedom) with inconsistent MDE predictions and noisy radar points leads to an overdetermined system (see Tab. 3). PolyRad offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
In terms of overhead, compared to predicting scale and shift coefficients for a linear transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost: an MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Supp. Mat).
Geometric Intuition. Applying polynomial fitting to the scaleless depth map can be viewed as a flexible, depth-dependent correction mechanism. In contrast to a single scale-and-shift operation, which uniformly adjusts the entire depth field by a single global factor, polynomial transformations introduce multiple degrees of freedom—higher-order terms enable correction of misalignments in relative depth between objects and local regions.
By introducing polynomial terms, our method enables depth adjustments that vary according to the initial depth estimates. Lower-order terms ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) capture broad scene structure, ensuring that near and far objects maintain reasonable depth relationships. Higher-order terms ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) refine depth predictions by introducing localized adjustments that correct subtle variations, such as mid-range depth discrepancies or fine object boundaries.
The concept of inflection points provides an intuitive understanding of this approach. Each additional polynomial order introduces more potential inflection points, allowing the depth transformation to shift curvature where needed. This flexibility enables “stretching” or “compressing” different depth levels: areas already near their correct metric depth receive minimal adjustment, while regions suffering larger errors (e.g., due to misalignment with other regions) undergo more substantial correction. Higher-order terms magnify small discrepancies in the initial MDE predictions, making them more apparent to the model and easier to correct. In this way, the degree of the polynomial, selected as a hyperparameter, directly reflects the capacity of our model to perform non-uniform depth-dependent adjustments.
The sign of the predicted polynomial coefficients furthers our interpretation. Positive coefficients for higher-order terms push depth values outward while negative ones pull them inward, effectively expanding or contracting selected depth intervals to correct local errors. In doing so, these coefficients shape the curvature of the polynomial, dictating where and how inflection points arise. For instance, if the model consistently overestimates distant objects, a negative high-order coefficient can compress that region, whereas a positive coefficient might be learned to correct underestimations. By treating coefficient signs as dynamic anchors for curvature changes, the polynomial fitting framework provides an interpretable mechanism (see Fig. 4) for how our method adjusts depth predictions at different scales.
Figure 4: PolyRad leverages higher-degree polynomials to reduce error by correcting depth misalignments using inflection points.
@@ 3.4 Loss Function
We employ a loss function comprising three terms to guide the learning of the polynomial coefficients and ensure accurate depth estimation. Our loss is defined as:
$\mathcal{L}=\lambda_{1}\|\hat{d}-d_{\text{gt}}\|_{1}+\lambda_{2}\|\hat{d}-d_{\text{gt}}\|_{2}^{2}+\lambda_{3}\left\|1-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (8)
@@
The novel component of our loss function lies in the third term, which encourages the first derivative of the predicted depth map $\hat{d}$ with respect to the input scaleless depth $z$ to generally be close to 1. Formally, we define:
@@
This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression [1]. Generally, within an object or local region, a pixel with a higher initial scaleless depth value should not be assigned a lower final metric depth value compared to a pixel with a lower initial scaleless depth. The preservation of ordinal relationships is an inductive bias grounded in the assumption that the MDE model provides a reasonably accurate estimation of intra-object relative depth.
However, polynomial fitting introduces a large function space, with expressiveness increasing as the polynomial degree $N$ grows, potentially disrupting local monotonic depth ordering if not properly constrained. Our regularization term addresses this by encouraging that, on average across the scene, the slope of our polynomial transformation remains close to 1, thus preserving monotonicity within local regions. At the same time, since there are fewer measurements between objects, PolyRad is free to correct cross-object misalignments while preserving local depth ordinality.
A first derivative close to 1 means that for every incremental change in the input scaleless depth $z$, there is an approximately equal incremental change in the predicted metric depth $\hat{d}$. This constraint ensures that our model does not overfit to noisy radar data or introduce spurious corrections by avoiding highly non-linear transformations that could disrupt the monotonicity of depth within local regions. Furthermore, this regularization term mitigates oscillatory polynomial curves, a known challenge when using higher-degree polynomials [3]. Without this constraint of regularity, the polynomial may excessively oscillate due to overfitting, leading to erratic depth predictions in some regions.

### v2

@@ 3.1 Motivation
indicating where the second derivative changes sign. By learning inflection points, we can model transition regions in the curvature of the MDE error, allowing necessary non-uniform corrections.
3.2 Representation Learning and Fusion
Radar Processing. We begin with a radar point cloud $C\in\mathbb{R}^{N_{C}\times 3}$, where each of the $N_{C}$ points is represented by its $(x,y,z)$ coordinates in three-dimensional camera space. Radar point clouds, while noisy and sparse, provide metric depth measurements. We concatenate a sinusoidal 3D positional embedding $\phi_{3D}(x,y,z)$ to $C$, and feed the resulting representation into a multilayer perceptron (MLP) $\psi_{r}$, producing radar features $F_{r}\in\mathbb{R}^{N_{C}\times c_{r}}$, where $c_{r}$ is the feature dimension.
Radar Aggregation. To facilitate effective radar feature aggregation, we introduce a set of learnable prototypes $P\in\mathbb{R}^{N_{P}\times c_{r}}$, where these $N_{P}$ prototypes capture diverse spatial and geometric properties present in the radar point cloud. Each prototype learns to focus on specific patterns in the configuration of radar points, making them highly expressive and adaptable across varying scenes. Unlike existing works (Singh et al., 2023; Li et al., 2024a) that directly encode and pool radar points—making them more susceptible to nuisances like multipath propagation—our learnable prototypes identify and match meaningful patterns within the point cloud. This allows for selective feature aggregation that mitigates the impact of outliers, resulting in more robust depth predictions.
Next, learnable prototypes are matched to the most relevant radar features for a given input. We treat the prototypes as centroids and perform soft clustering over the radar features. Each radar point is softly assigned to prototypes according to feature similarity, and the corresponding values are aggregated to yield a global scene-descriptive representation $F_{R}$:
$D_{ij}=\|P_{j}-\Phi_{r}(F_{r})_{i}\|^{2},\quad F_{R}=\text{softmax}\!\left(-D\,/\,\tau\right)\,\Psi_{r}(F_{r}),$ (2)
where $\Phi_{r}$, $\Psi_{r}$ are MLPs that project the radar features $F_{r}$, the matrix $D\in\mathbb{R}^{N_{P}\times N_{C}}$ contains the pairwise squared distances between prototypes $\{P_{j}\}_{j=1}^{N_{P}}$ and projected radar features $\{\Phi_{r}(F_{r})_{i}\}_{i=1}^{N_{C}}$, and $\tau$ denotes temperature. This clustering-based formulation enables prototypes to capture recurring spatial and geometric patterns in radar point clouds. Comparatively, applying lidar depth completion methods, which densify a sparse projection of points using surrounding context, to a radar point cloud results in poor performance (see Sec. E). This is because radar measurements are orders of magnitude sparser than lidar and are much noisier, especially when lacking sufficient antenna elements, e.g., elevation ambiguity, or range resolution (Singh et al., 2023).
MDE-Radar Fusion. The scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ from the MDE model is encoded with a learnable depth encoder $f_{z}$, producing depth features $Z\in\mathbb{R}^{(H\times W)\times c_{z}}$. These features inherit invariants learned through large-scale MDE training—such as robustness to color variations, illumination changes, and diverse object poses—and therefore primarily encode object-level properties. Since variations in the shape of objects and their geometry are generally more stable across scenes than photometric appearance, the resulting depth features provide a more reliable geometric context for fusion with radar features. This allows radar point configurations to be matched to consistently observed shapes and surfaces, rather than to pixel-wise color intensities that can vary arbitrarily with lighting and viewpoint. To this end, we learn soft spatial correspondences between the depth features $Z$ and the radar features $F_{R}$ to construct a unified scene representation $S$ that fuses the structural information encoded in the MDE predictions with the metric cues provided by the radar features:
$S=\text{softmax}\left(\frac{(Z+E)\times(\Phi_{R}(F_{R}))^{T}}{\sqrt{c_{r}}}\right)\Psi_{R}(F_{R}),$ (3)
where $E\in\mathbb{R}^{(H\times W)\times c_{z}}$ is a learned 2D positional embedding, and $\Phi_{R}$, $\Psi_{R}$ are MLPs that project the aggregated radar features $F_{R}$. $c_{z}$ is set equal to $c_{r}$ to align the depth and radar features within a common embedding space, facilitating cross-modal fusion, and $S$ is reshaped to be in $\mathbb{R}^{H\times W\times c_{z}}$.
Predicting Coefficients. The fused representation $S$ is passed through a shallow convolutional neural network (CNN) $f_{s}$ followed by a global average pooling (GAP) layer to yield $\bar{S}=\text{GAP}(f_{s}(S))$, which is a $c_{s}$-dimensional feature vector. The GAP layer aggregates spatial information from the entire scene, ensuring that $\bar{S}$ captures global context. This final scene representation $\bar{S}$ is then fed into an MLP $\psi_{s}$, which predicts the $N+1$ polynomial coefficients as a vector $\hat{c}$:
$\hat{c}=\psi_{s}(\bar{S})\in\mathbb{R}^{N+1}.$ (4)
These coefficients allow the model to adaptively refine the initial depth predictions from the MDE model, with each coefficient adjusting depth at different scales and granularities. Lower-order terms capture broad scene structure, while higher-order terms enable fine-grained and cross-region corrections, resulting in a high-fidelity metric-scale depth map.
3.3 Polynomial Fitting
Given a scaleless depth map $z\in\mathbb{R}_{+}^{H\times W}$ predicted by the MDE foundation model, our goal is to transform it into a metric depth map $\hat{d}\in\mathbb{R}_{+}^{H\times W}$ using the polynomial coefficients $\{\hat{c}_{0},\hat{c}_{1},\dots,\hat{c}_{N}\}$ predicted by our network (Sec. 3.2). Formally, we express the final depth map $\hat{d}$ as:
$\begin{split}\hat{d}(x,y)=\sum_{i=0}^{N}\,\hat{c}_{i}\cdot z(x,y)^{i},\;\;\forall(x,y)\in H\times W.\end{split}$ (5)
Figure 3: POLAR leverages spatial information from radar points to predict higher-degree polynomial transformations that can correct non-affine errors in MDE predictions.
This polynomial formulation applies the learned coefficients to successive powers of the scaleless depth map, enabling complex, non-linear transformations that surpass the limitations of affine transformations. Each pixel’s final depth is thus computed by summing multiple weighted terms, where each term corresponds to a different polynomial order of the initial depth prediction.
This is in contrast to directly mapping to a metric-scale depth map (Singh et al., 2023; Li et al., 2024a), where the degrees of freedom comprise every predicted pixel. Each corresponding point and pixel prediction can be viewed as a constraint or equation in our estimation system. As the radar point cloud ($\sim 10^{2}$ points) only occupies a small subset of the image space ($\sim 10^{6}$ pixels), the solution is underdetermined and often irregular (see Fig. 4). Conversely, a simple linear fit (only two degrees of freedom) with inconsistent MDE predictions and noisy radar points leads to an overdetermined system (see Tab. 6). POLAR offers the middle ground: By a flexible choice of polynomial degree, we can appropriately tune and select the complexity of the function, i.e., degrees of freedom, that leads to a better fit between MDE predictions and radar points.
In terms of overhead, compared to predicting two coefficients (scale and shift) for an affine transformation, predicting $N+1$ coefficients for polynomial fitting introduces a negligible increase in computational cost ($0.01\%$ more FLOPs): the final MLP outputs an $(N+1)$-dimensional vector instead of a two-dimensional vector, and the final depth map is computed through parallelized exponentiation and multiplication operations (see Sec. B for the complete derivation).
Geometric Intuition. Applying polynomial fitting to the scaleless depth map is best viewed as a flexible, depth-dependent correction mechanism. In contrast to a single scale-and-shift operation, which uniformly adjusts the entire depth field by a single global factor, polynomial transformations introduce multiple degrees of freedom—higher-order terms enable correction of misalignments in relative depth between objects and local regions.
By introducing polynomial terms, our method enables depth adjustments that vary according to the initial depth estimates $z$. Lower-order coefficients ($\hat{c}_{0}$, $\hat{c}_{1}$, $\cdots$) establish a broad global scale, providing a metric basis for the initially scaleless $z$. Higher-order coefficients ($\cdots$, $\hat{c}_{N-1}$, $\hat{c}_{N}$) introduce both low-frequency corrections, e.g., resolving cross-object misalignments (Fig. 1), as well as high-frequency corrections, e.g., sharpening fine object boundaries (Fig. 4, Image C).
The concept of inflection points provides an intuitive understanding of this approach. Each additional polynomial order introduces more potential inflection points, allowing the depth transformation to shift curvature where needed. This flexibility enables “stretching” or “compressing” different depth levels: areas already near their correct metric depth receive minimal adjustment, while regions suffering larger errors (e.g., due to misalignment with other regions) undergo more substantial correction. Higher-order terms magnify small (i.e., high-frequency) discrepancies in the initial MDE predictions, making them more apparent to the model and easier to correct. In this way, the degree of the polynomial, selected as a hyperparameter, governs the capacity of our model to apply non-uniform corrections across the scene.
The sign of the predicted polynomial coefficients furthers our interpretation. Positive coefficients for higher-order terms push depth values outward while negative ones pull them inward, effectively expanding or contracting selected depth intervals to correct local errors. In doing so, these coefficients shape the curvature of the polynomial, dictating where and how inflection points arise. For instance, if the model consistently overestimates distant objects, a negative high-order coefficient can compress that region, whereas a positive coefficient might be learned to correct underestimations. By treating coefficient signs as dynamic anchors for curvature changes, the polynomial fitting framework provides an interpretable mechanism (e.g., Fig. 3) for refining scaleless depth into high-fidelity metric depth.
Table 1: Quantitative results on the nuScenes, ZJU-4DRadarCam (ZJU), and View-of-Delft test sets evaluated with various maximum evaluation distances.
@@ 3.3 Polynomial Fitting
   50m RadarNet [CVPR ’23] 1727.7 3746.8 1430.5 3250.8 2944.8 6113.2
SparseBeatsDense [ECCV ’24] 1524.5 3567.3 1424.4 3267.5 2909.9 5746.4
GET-UP [WACV ’25] 1241.0 2857.0 1483.9 3220.5 2330.0 4565.0
RadarCam-Depth [ICRA ’24] 1286.1 2964.3 1067.5 2817.4 1895.1 4458.7
TacoDepth [CVPR ’25] 1046.8 2487.5 930.2 2477.3 - -
POLAR (Ours) 1014.4 2475.7 578.0 1108.6 1293.7 2420.6
   70m RadarNet [CVPR ’23] 2073.2 4590.7 1543.8 3655.3 3428.9 7331.9
SparseBeatsDense [ECCV ’24] 1822.9 4303.6 1520.0 3593.4 3408.3 6914.7
GET-UP [WACV ’25] 1541.0 3657.0 1651.5 3711.7 2758.0 5678.0
RadarCam-Depth [ICRA ’24] 1587.9 3662.5 1157.0 3117.7 2095.4 4944.6
TacoDepth [CVPR ’25] 1347.1 3152.8 983.1 2779.6 - -
POLAR (Ours) 1286.1 2947.3 603.7 1154.9 1442.9 2803.8
   80m RadarNet [CVPR ’23] 2179.3 4898.7 1578.4 3804.2 3597.0 7809.2
SparseBeatsDense [ECCV ’24] 1927.0 4609.6 1548.4 3708.1 3584.0 7375.2
GET-UP [WACV ’25] 1632.0 3932.0 1699.7 3882.6 2917.3 6145.1
RadarCam-Depth [ICRA ’24] 1689.7 3948.0 1183.5 3229.0 2227.4 5385.8
TacoDepth [CVPR ’25] 1492.4 3324.8 1032.5 2850.3 - -
POLAR (Ours) 1407.8 3193.5 629.6 1171.3 1500.1 3951.8
@@ 3.4 Loss Function
We employ a loss function comprising three terms, each weighted by its $\lambda$, to guide the learning of the polynomial coefficients and ensure accurate depth estimation. Our loss is defined as:
$\mathcal{L}=\lambda_{L_{1}}\|\hat{d}-d_{\text{gt}}\|_{1}+\lambda_{L_{2}}\|\hat{d}-d_{\text{gt}}\|_{2}^{2}+\lambda_{\text{mono}}\left\|\mathbf{1}_{H\times W}-\frac{\mathrm{d}\hat{d}}{\mathrm{d}z}\right\|_{1},$ (6)
@@
The novel component of our objective lies in the third term, which constrains the first derivative of the predicted depth $\hat{d}$ w.r.t the input scaleless depth $z$ (Eq. 7) to remain near that of $\hat{d}=z$.
@@
This regularization enforces that our polynomial fitting function remains approximately monotonically increasing, akin to isotonic regression (Barlow et al., 1972). Generally, within an object or local region, a pixel with a higher initial scaleless depth value should not be assigned a lower final metric depth value compared to a pixel with a lower initial scaleless depth. The preservation of ordinal relationships is an inductive bias grounded in the assumption that the MDE model provides a reasonably accurate estimation of intra-object relative depth.
Polynomial fitting introduces a large function space, with expressiveness growing with degree $N$, which can significantly and detrimentally disrupt local monotonic depth ordering if unconstrained. Our regularization term addresses this by preserving local depth ordinality while still allowing corrections of cross-object misalignments. This constraint prevents overfitting of nonlinear transformations to noisy radar data, spurious correlations, and outliers, thereby avoiding the oscillatory behavior that is a known challenge with fitting higher-degree polynomials (Bishop, 2006).
Figure 4: Qualitative results on nuScenes. GET-UP and RadarCam-Depth (RC-D) fail to reconstruct entire regions, yielding objects with large depth errors. Raw MDE yields reasonable relative reconstructions but suffers from incorrect global scale and cross-object misalignments. POLAR leverages polynomial fitting to recover a global scale and correct these misalignments. See Fig. 5 for colorbars.